# Light-induced ultrafast magnetization dynamics in van der Waals antiferromagnetic CrSBr

### v2

@@ Acknowledgements.
This research was supported by the US National Science Foundation (NSF) through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824. The supercomputing time was provided by DARWIN (Delaware Advanced Research Workforce and Innovation Network), which is supported by NSF Grant No. MRI-1919839.References
@@
- Kuo et al. [2016] C.-T. Kuo, M. Neumann, K. Balamurugan, H. J. Park, S. Kang, H. W. Shiu, J. H. Kang, B. H. Hong, M. Han, T. W. Noh, et al., Exfoliation and raman spectroscopic fingerprint of few-layer nips3 van der waals crystals, Scientific reports 6, 20904 (2016).
- Lin et al. [2016] M.-W. Lin, H. L. Zhuang, J. Yan, T. Z. Ward, A. A. Puretzky, C. M. Rouleau, Z. Gai, L. Liang, V. Meunier, B. G. Sumpter, et al., Ultrathin nanosheets of crsite 3: a semiconducting two-dimensional ferromagnetic material, Journal of Materials Chemistry C 4, 315 (2016).
- Lee et al. [2016] S. Lee, K.-Y. Choi, S. Lee, B. H. Park, and J.-G. Park, Tunneling transport of mono-and few-layers magnetic van der waals mnps3, APL Materials 4, 1 (2016).
@@
- Wu et al. [2019] M. Wu, Z. Li, T. Cao, and S. G. Louie, Physical origin of giant excitonic and magneto-optical responses in two-dimensional ferromagnetic insulators, Nature communications 10, 2371 (2019).
- Brennan et al. [2024] N. J. Brennan, C. A. Noble, J. Tang, M. E. Ziebel, and Y. J. Bae, Important elements of spin-exciton and magnon-exciton coupling, ACS Physical Chemistry Au 4, 322 (2024).
- Ahn et al. [2024] Y. Ahn, X. Guo, S. Son, Z. Sun, and L. Zhao, Progress and prospects in two-dimensional magnetism of van der waals materials, Progress in Quantum Electronics 93, 100498 (2024).
- Song and Xu [2024] T. Song and X. Xu, The future of 2d spintronics, Nature Reviews Electrical Engineering 1, 696 (2024).
- Geim and Grigorieva [2013] A. K. Geim and I. V. Grigorieva, Van der waals heterostructures, Nature 499, 419 (2013).
- Huang et al. [2018] B. Huang, G. Clark, D. R. Klein, D. MacNeill, E. Navarro-Moratalla, K. L. Seyler, N. Wilson, M. A. McGuire, D. H. Cobden, D. Xiao, et al., Electrical control of 2d magnetism in bilayer cri3, Nature nanotechnology 13, 544 (2018).
- Jiang et al. [2018] S. Jiang, J. Shan, and K. F. Mak, Electric-field switching of two-dimensional van der waals magnets, Nature materials 17, 406 (2018).
- Zhang et al. [2020] X.-X. Zhang, L. Li, D. Weber, J. Goldberger, K. F. Mak, and J. Shan, Gate-tunable spin waves in antiferromagnetic atomic bilayers, Nature materials 19, 838 (2020).
@@
- Diederich et al. [2023] G. M. Diederich, J. Cenker, Y. Ren, J. Fonseca, D. G. Chica, Y. J. Bae, X. Zhu, X. Roy, T. Cao, D. Xiao, et al., Tunable interaction between excitons and hybridized magnons in a layered semiconductor, Nature Nanotechnology 18, 23 (2023).
- Zhang et al. [2022] P. Zhang, T.-F. Chung, Q. Li, S. Wang, Q. Wang, W. L. Huey, S. Yang, J. E. Goldberger, J. Yao, and X. Zhang, All-optical switching of magnetization in atomically thin cri3, Nature materials 21, 1373 (2022).
- Hu et al. [2022] H. Hu, N. Chen, H. Teng, R. Yu, Y. Qu, J. Sun, M. Xue, D. Hu, B. Wu, C. Li, et al., Doping-driven topological polaritons in graphene/-moo3 heterostructures, Nature Nanotechnology 17, 940 (2022).
- Ni et al. [2021] Z. Ni, A. Haglund, H. Wang, B. Xu, C. Bernhard, D. Mandrus, X. Qian, E. Mele, C. Kane, and L. Wu, Imaging the néel vector switching in the monolayer antiferromagnet mnpse3 with strain-controlled ising order, Nature nanotechnology 16, 782 (2021).
@@
- Klaproth et al. [2023] T. Klaproth, S. Aswartham, Y. Shemerliuk, S. Selter, O. Janson, J. van den Brink, B. Büchner, M. Knupfer, S. Pazek, D. Mikhailova, et al., Origin of the magnetic exciton in the van der waals antiferromagnet nips3, Physical Review Letters 131, 256504 (2023).
- Ziebel et al. [2024] M. E. Ziebel, M. L. Feuer, J. Cox, X. Zhu, C. R. Dean, and X. Roy, Crsbr: an air-stable, two-dimensional magnetic semiconductor, Nano Letters 24, 4319 (2024).
- Lauk et al. [2020] N. Lauk, N. Sinclair, S. Barzanjeh, J. P. Covey, M. Saffman, M. Spiropulu, and C. Simon, Perspectives on quantum transduction, Quantum Science and Technology 5, 020501 (2020).
- Liu et al. [2023] S. Liu, I. A. Malik, V. L. Zhang, and T. Yu, Lightning the spin: Harnessing the potential of 2d magnets in opto-spintronics, Advanced Materials , 2306920 (2023).
- Gish et al. [2024] J. T. Gish, D. Lebedev, T. W. Song, V. K. Sangwan, and M. C. Hersam, Van der waals opto-spintronics, Nature Electronics 7, 336 (2024).
- Kirilyuk et al. [2010] A. Kirilyuk, A. V. Kimel, and T. Rasing, Ultrafast optical manipulation of magnetic order, Reviews of Modern Physics 82, 2731 (2010).
- Basov et al. [2017] D. Basov, R. Averitt, and D. Hsieh, Towards properties on demand in quantum materials, Nature materials 16, 1077 (2017).
- De La Torre et al. [2021] A. De La Torre, D. M. Kennes, M. Claassen, S. Gerber, J. W. McIver, and M. A. Sentef, Colloquium: Nonthermal pathways to ultrafast control in quantum materials, Reviews of Modern Physics 93, 041002 (2021).
- Beaurepaire et al. [1966] E. Beaurepaire, J.-C. Merle, A. Daunois, and J.-Y. Bigot, Ultrafast spin dynamics in ferromagnetic nickel, Physical review letters 76, 4250 (1966).
- Dirnberger et al. [2023] F. Dirnberger, J. Quan, R. Bushati, G. M. Diederich, M. Florian, J. Klein, K. Mosina, Z. Sofer, X. Xu, A. Kamra, et al., Magneto-optics in a van der waals magnet tuned by self-hybridized polaritons, Nature 620, 533 (2023).
- Sun et al. [2024] Y. Sun, F. Meng, C. Lee, A. Soll, H. Zhang, R. Ramesh, J. Yao, Z. Sofer, and J. Orenstein, Dipolar spin wave packet transport in a van der waals antiferromagnet, Nature Physics 20, 794–800 (2024).
- Diederich et al. [2025] G. M. Diederich, M. Nguyen, J. Cenker, J. Fonseca, S. Pumulo, Y. J. Bae, D. G. Chica, X. Roy, X. Zhu, D. Xiao, et al., Exciton dressing by extreme nonlinear magnons in a layered semiconductor, Nature Nanotechnology 20, 1 (2025).
@@
- Yang et al. [2021] K. Yang, G. Wang, L. Liu, D. Lu, and H. Wu, Triaxial magnetic anisotropy in the two-dimensional ferromagnetic semiconductor crsbr, Physical Review B 104, 144416 (2021).
@@
- López-Paz et al. [2022] S. A. López-Paz, Z. Guguchia, V. Y. Pomjakushin, C. Witteveen, A. Cervellino, H. Luetkens, N. Casati, A. F. Morpurgo, and F. O. von Rohr, Dynamic magnetic crossover at the origin of the hidden-order in van der waals antiferromagnet crsbr, Nature Communications 13, 4745 (2022).
- Elliott et al. [2020] P. Elliott, N. Singh, K. Krieger, E. Gross, S. Sharma, and J. Dewhurst, The microscopic origin of spin-orbit mediated spin-flips, Journal of Magnetism and Magnetic Materials 502, 166473 (2020).
- Thorpe et al. [2023] A. Thorpe, N. Boroumand, A. Parks, E. Goulielmakis, and T. Brabec, High harmonic generation in solids: Real versus virtual transition channels, Physical Review B 107, 075135 (2023).
- Zhang et al. [2018] G. Zhang, M. Si, M. Murakami, Y. Bai, and T. F. George, Generating high-order optical and spin harmonics from ferromagnetic monolayers, Nature Communications 9, 3031 (2018).

### v3

@@ Acknowledgements.
Acknowledgments— This research was supported by the US National Science Foundation (NSF) through the University of Delaware Materials Research Science and Engineering Center, DMR-2011824, and the University of Delaware (UD) Research Foundation Strategic Initiative Award. The supercomputing time was provided by DARWIN (Delaware Advanced Research Workforce and Innovation Network), which is supported by NSF Grant No. MRI-1919839.References
- Žutić et al. [2004] I. Žutić, J. Fabian, and S. D. Sarma, Spintronics: Fundamentals and applications, Rev. Mod. Phys. 76, 323 (2004).
- Baltz et al. [2018] V. Baltz, A. Manchon, M. Tsoi, T. Moriyama, T. Ono, and Y. Tserkovnyak, Antiferromagnetic spintronics, Rev. Mod. Phys. 90, 015005 (2018).
- Šmejkal et al. [2018] L. Šmejkal, Y. Mokrousov, B. Yan, and A. H. MacDonald, Topological antiferromagnetic spintronics, Nat. Phys. 14, 242 (2018).
- Pirro et al. [2021] P. Pirro, V. I. Vasyuchka, A. A. Serga, and B. Hillebrands, Advances in coherent magnonics, Nature Reviews Materials 6, 1114 (2021).
- Kirilyuk et al. [2010] A. Kirilyuk, A. V. Kimel, and T. Rasing, Ultrafast optical manipulation of magnetic order, Rev. Mod. Phys. 82, 2731 (2010).
- Beaurepaire et al. [1966] E. Beaurepaire, J.-C. Merle, A. Daunois, and J.-Y. Bigot, Ultrafast spin dynamics in ferromagnetic nickel, Phys. Rev. Lett. 76, 4250 (1966).
- Bigot et al. [2009] J.-Y. Bigot, M. Vomir, and E. Beaurepaire, Coherent ultrafast magnetism induced by femtosecond laser pulses, Nat. Phys. 5, 515 (2009).
- Siegrist et al. [2019] F. Siegrist, J. A. Gessner, M. Ossiander, C. Denker, Y.-P. Chang, M. C. Schröder, A. Guggenmos, Y. Cui, J. Walowski, U. Martens, et al., Light-wave dynamic control of magnetism, Nature 571, 240 (2019).
- Kholid et al. [2023] F. N. Kholid, D. Hamara, A. F. B. Hamdan, G. Nava Antonio, R. Bowen, D. Petit, R. Cowburn, R. V. Pisarev, D. Bossini, J. Barker, et al., The importance of the interface for picosecond spin pumping in antiferromagnet-heavy metal heterostructures, Nat. Commun. 14, 538 (2023).
- Golias et al. [2021] E. Golias, I. Kumberg, I. Gelen, S. Thakur, J. Gördes, R. Hosseinifar, Q. Guillet, J. Dewhurst, S. Sharma, C. Schüßler-Langeheine, et al., Ultrafast optically induced ferromagnetic state in an elemental antiferromagnet, Phys. Rev. Lett. 126, 107202 (2021).
- Seifert et al. [2016] T. Seifert, S. Jaiswal, U. Martens, J. Hannegan, L. Braun, P. Maldonado, F. Freimuth, A. Kronenberg, J. Henrizi, I. Radu, et al., Efficient metallic spintronic emitters of ultrabroadband terahertz radiation, Nat. Photonics. 10, 483 (2016).
- Wu et al. [2021] W. Wu, C. Yaw Ameyaw, M. F. Doty, and M. B. Jungfleisch, Principles of spintronic THz emitters, J. Appl. Phys. 130, 091101 (2021).
@@
- Kuo et al. [2016] C.-T. Kuo, M. Neumann, K. Balamurugan, H. J. Park, S. Kang, H. W. Shiu, J. H. Kang, B. H. Hong, M. Han, T. W. Noh, et al., Exfoliation and raman spectroscopic fingerprint of few-layer NiPS3 van der Waals crystals, Sci. Rep. 6, 20904 (2016).
- Lin et al. [2016] M.-W. Lin, H. L. Zhuang, J. Yan, T. Z. Ward, A. A. Puretzky, C. M. Rouleau, Z. Gai, L. Liang, V. Meunier, B. G. Sumpter, et al., Ultrathin nanosheets of CrSiTe3: a semiconducting two-dimensional ferromagnetic material, J. Mater. Chem. C. 4, 315 (2016).
- Lee et al. [2016] S. Lee, K.-Y. Choi, S. Lee, B. H. Park, and J.-G. Park, Tunneling transport of mono-and few-layers magnetic van der Waals MnPS3, APL Mater. 4, 1 (2016).
@@
- Wu et al. [2019] M. Wu, Z. Li, T. Cao, and S. G. Louie, Physical origin of giant excitonic and magneto-optical responses in two-dimensional ferromagnetic insulators, Nat. Commun. 10, 2371 (2019).
- Brennan et al. [2024] N. J. Brennan, C. A. Noble, J. Tang, M. E. Ziebel, and Y. J. Bae, Important elements of spin-exciton and magnon-exciton coupling, ACS Phys. Chem. Au. 4, 322 (2024).
- Ahn et al. [2024] Y. Ahn, X. Guo, S. Son, Z. Sun, and L. Zhao, Progress and prospects in two-dimensional magnetism of van der Waals materials, Prog Quantum Electron. 93, 100498 (2024).
- Song and Xu [2024] T. Song and X. Xu, The future of 2D spintronics, Nat. Rev. Electr. Eng. 1, 696 (2024).
- Da̧browski et al. [2022] M. Dąbrowski, M., S. Guo, M. Strungaru, P. S. Keatley, F. Withers, E. J. Santos, and R. J. Hicken, All-optical control of spin in a 2D van der Waals magnet, Nat. Commun. 13, 5976 (2022).
- Wu et al. [2024] D. Wu, M. Ye, H. Chen, Y. Xu, and W. Duan, Giant and controllable nonlinear magneto-optical effects in two-dimensional magnets, npj Comput. Mater. 10, 79 (2024).
- Gish et al. [2024] J. T. Gish, D. Lebedev, T. W. Song, V. K. Sangwan, and M. C. Hersam, Van der Waals opto-spintronics, Nat. Electron. 7, 336 (2024).
- Zhang et al. [2022] P. Zhang, T.-F. Chung, Q. Li, S. Wang, Q. Wang, W. L. Huey, S. Yang, J. E. Goldberger, J. Yao, and X. Zhang, All-optical switching of magnetization in atomically thin CrI3, Nat. Mater. 21, 1373 (2022).
- Huang et al. [2018] B. Huang, G. Clark, D. R. Klein, D. MacNeill, E. Navarro-Moratalla, K. L. Seyler, N. Wilson, M. A. McGuire, D. H. Cobden, D. Xiao, et al., Electrical control of 2D magnetism in bilayer CrI3, Nat. Nanotechnol. 13, 544 (2018).
- Jiang et al. [2018] S. Jiang, J. Shan, and K. F. Mak, Electric-field switching of two-dimensional van der Waals magnets, Nat. Mater. 17, 406 (2018).
- Zhang et al. [2020] X.-X. Zhang, L. Li, D. Weber, J. Goldberger, K. F. Mak, and J. Shan, Gate-tunable spin waves in antiferromagnetic atomic bilayers, Nat. Mater. 19, 838 (2020).
@@
- Diederich et al. [2023] G. M. Diederich, J. Cenker, Y. Ren, J. Fonseca, D. G. Chica, Y. J. Bae, X. Zhu, X. Roy, T. Cao, D. Xiao, et al., Tunable interaction between excitons and hybridized magnons in a layered semiconductor, Nat. Nanotechnol. 18, 23 (2023).
- Hu et al. [2022] H. Hu, N. Chen, H. Teng, R. Yu, Y. Qu, J. Sun, M. Xue, D. Hu, B. Wu, C. Li, et al., Doping-driven topological polaritons in graphene/-MoO3 heterostructures, Nat. Nanotechnol. 17, 940 (2022).
- Ni et al. [2021] Z. Ni, A. Haglund, H. Wang, B. Xu, C. Bernhard, D. Mandrus, X. Qian, E. Mele, C. Kane, and L. Wu, Imaging the Néel vector switching in the monolayer antiferromagnet MnPSe3 with strain-controlled Ising order, Nat. Nanotechnol. 16, 782 (2021).
- Diederich et al. [2025] G. M. Diederich, M. Nguyen, J. Cenker, J. Fonseca, S. Pumulo, Y. J. Bae, D. G. Chica, X. Roy, X. Zhu, D. Xiao, et al., Exciton dressing by extreme nonlinear magnons in a layered semiconductor, Nat. Nanotechnol. 20, 1 (2025).
@@
- Klaproth et al. [2023] T. Klaproth, S. Aswartham, Y. Shemerliuk, S. Selter, O. Janson, J. van den Brink, B. Büchner, M. Knupfer, S. Pazek, D. Mikhailova, et al., Origin of the magnetic exciton in the van der Waals antiferromagnet NiPS3, Phys. Rev. Lett. 131, 256504 (2023).
- Varela-Manjarres et al. [2025] J. Varela-Manjarres, Y. Ren, and B. K. Nikolic, Ultrafast optical excitation of magnons in 2D antiferromagnets via spin torque exerted by photocurrent of excitons: Signatures in charge pumping and THz emission, arXiv:2507.02793 (2025).
- Chen et al. [2019] Z. Chen, J.-W. Luo, and L.-W. Wang, Revealing angular momentum transfer channels and timescales in the ultrafast demagnetization process of ferromagnetic semiconductors, PNAS 116, 19258 (2019).
- He et al. [2021] J. He, S. Li, A. Bandyopadhyay, and T. Frauenheim, Unravelling photoinduced interlayer spin transfer dynamics in two-dimensional nonmagnetic-ferromagnetic van der Waals heterostructures, Nano Lett. 21, 3237 (2021).
- Li et al. [2022] S. Li, L. Zhou, T. Frauenheim, and J. He, Light-controlled ultrafast magnetic state transition in antiferromagnetic–ferromagnetic van der Waals heterostructures, J. Phys. Chem. Lett. 13, 6223 (2022).
- He et al. [2023] J. He, S. Li, T. Frauenheim, and Z. Zhou, Ultrafast laser pulse induced transient ferrimagnetic state and spin relaxation dynamics in two-dimensional antiferromagnets, Nano Lett. 23, 8348 (2023).
- Zhou et al. [2023] Z. Zhou, Z. Zheng, J. He, J. Wang, O. V. Prezhdo, and T. Frauenheim, Ultrafast laser control of antiferromagnetic–ferrimagnetic switching in two-dimensional ferromagnetic semiconductor heterostructures, Nano Lett. 23, 5688 (2023).
- Guo et al. [2024] Y. Guo, Y. Zhang, Q. L. Liu, Z. Zhou, J. He, S. Yuan, T. Heine, and J. Wang, Laser-induced ultrafast spin injection in all-semiconductor magnetic CrI3/WSe2 heterobilayer, ACS Nano 18, 11732 (2024).
@@
- Ziebel et al. [2024] M. E. Ziebel, M. L. Feuer, J. Cox, X. Zhu, C. R. Dean, and X. Roy, CrSBr: an air-stable, two-dimensional magnetic semiconductor, Nano Lett. 24, 4319 (2024).
- Tengdin et al. [2018] P. Tengdin, W. You, C. Chen, X. Shi, D. Zusin, Y. Zhang, C. Gentry, A. Blonsky, M. Keller, P. M. Oppeneer, et al., Critical behavior within 20 fs drives the out-of-equilibrium laser-induced magnetic phase transition in nickel, Sci. Adv. 4, eaap9744 (2018).
- Krieger et al. [2015] K. Krieger, J. K. Dewhurst, P. Elliott, S. Sharma, and E. K. U. Gross, Laser-induced demagnetization at ultrashort time scales: Predictions of TDDFT, J. Chem. Theory Comput. 11, 4870 (2015).
- Elliott et al. [2020] P. Elliott, N. Singh, K. Krieger, E. Gross, S. Sharma, and J. Dewhurst, The microscopic origin of spin-orbit mediated spin-flips, J. Magn. Magn. Mater. 502, 166473 (2020).
@@
- Yang et al. [2021] K. Yang, G. Wang, L. Liu, D. Lu, and H. Wu, Triaxial magnetic anisotropy in the two-dimensional ferromagnetic semiconductor CrSBr, Phys. Rev. B. 104, 144416 (2021).
@@
- López-Paz et al. [2022] S. A. López-Paz, Z. Guguchia, V. Y. Pomjakushin, C. Witteveen, A. Cervellino, H. Luetkens, N. Casati, A. F. Morpurgo, and F. O. von Rohr, Dynamic magnetic crossover at the origin of the hidden-order in van der Waals antiferromagnet CrSBr, Nat. Commun. 13, 4745 (2022).
- [56] The Supplemental Material includes the band structure, partial density of states, charge currents, magnetization dynamics at high fluence case with above gap frequency, influence of magnetic fields, spin and layer resolved charge dynamics and high-harmonic generation. Refs. 54, 61 are cited.
- Wilson et al. [2021] N. P. Wilson, K. Lee, J. Cenker, K. Xie, A. H. Dismukes, E. J. Telford, J. Fonseca, S. Sivakumar, C. Dean, T. Cao, et al., Interlayer electronic coupling on demand in a 2D magnetic semiconductor, Nat. Mater. 20, 1657 (2021).
- Ruhman et al. [1988] S. Ruhman, A. G. Joly, and K. A. Nelson, Coherent molecular vibrational motion observed in the time domain through impulsive stimulated raman scattering, IEEE Journal of Quantum Electronics 24, 460 (1988).
- Weiner et al. [1991] A. M. Weiner, D. Leaird, G. P. Wiederrecht, and K. A. Nelson, Femtosecond multiple-pulse impulsive stimulated raman scattering spectroscopy, Journal of the Optical Society of America B 8, 1264 (1991).
- Zeiger et al. [1992] H. Zeiger, J. Vidal, T. Cheng, E. Ippen, G. Dresselhaus, and M. Dresselhaus, Theory for displacive excitation of coherent phonons, Phys. Rev. B. 45, 768 (1992).
- Zhang et al. [2018] G. Zhang, M. Si, M. Murakami, Y. Bai, and T. F. George, Generating high-order optical and spin harmonics from ferromagnetic monolayers, Nat. Commun. 9, 3031 (2018).